[figure: 9.7 cm]

Neutron diffraction study of crystal and magnetic structure of (La_{0.7}Ca_{0.3})_{1−x}Mn_{1+x}O_{3} (x = 0, 0.1) compounds

W. Bażela^{a}, V. Dyakonov^{b,c}, V. Pashchenko^{c}, B. Penc^{d}, H. Szymczak^{b}, J.H.

Hernández-Velasco^{e} and A. Szytuła^{d} a Institute of Physics, Technical University of Cracow, 30-084 Kraków, Poland b Institute of Physics, Polish Academy of Sciences, 02-668 Warszawa, Poland c A.D. Galkin Donetsk Physico-Technical Institute, 340114 Donetsk, Ukraine d Institute of Physics, Jagiellonian University, 30-059 Kraków, Poland e BENSC, Hahn-Meitner Institute , D-14 109 Berlin-Wannsee, Germany

Evolution of crystal and magnetic structures of (La_{0.7}Ca_{0.3})_{1−x}Mn_{1+x}O_{3} (x = 0, 0.1) was determined as a function of temperature. For both compounds the crystal structure is orthorhombic (space group Pnma) with the following sequence of the lattice constants b/\sqrt{2} < a < c for x=0 and c < a < b/\sqrt{2} for x=0.1. The para-ferromagnetic transition at T 260 K is accompanied by structural distortions. For x=0 the ferromagnetic ordering of the Mn moments equal 3.6(1) μ_{B}, is stable in the temperature range from 1.5 K up to T_{c}=265 K. For x=0.1 at 1.5 K the magnetic order is complex, with the ferromagnetic (μ=3.5(1) μ_{B}) and modulated (0.5(1) μ_{B}) component. This order is stable up to 45 K. Then up to T_{c}=260 K the ferromagnetic order exists. The obtained results correlate well with the macroscopic magnetic data [1]. Reference.

[1] V. Dyakonov et al., J. Magn. Magn. Mater. (in press).

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : A. Szytuła

Address for correspondence : Institute of Physics, Jagiellonian University, Reymonta 4, 30-059 Kraków, Poland

Email address : szytula@if.uj.edu.pl